# A geometric perspective on Algebraic Quantum Field Theory

### v1

@@ 2.2 Euler elements
We now introduce the natural general setting, where all these examples naturally fit in and the previous difference about the inclusion properties of wedge regions becomes clear. The primary reference for this is [MN21].
@@
Then $\tau_{h}(y_{j})=\exp(i\pi\mathop{{\rm ad}}\nolimits h)(y_{j})=(-1)^{j}y_{j}$ for $y_{j}\in{\mathfrak{g}}_{j}(h)$ defines an involutive automorphism of ${\mathfrak{g}}$ called an Euler involution.
We denote $\mathcal{E}({\mathfrak{g}})$ the set of Euler elements in ${\mathfrak{g}}$. The orbit of an Euler element $h$ under the group $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})=\langle e^{\mathop{{\rm ad}}\nolimits{\mathfrak{g}}}\rangle$ of inner automorphisms is denoted with $\mathcal{O}_{h}=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})h\subseteq{\mathfrak{g}}$. We say that $h$ is symmetric if $-h\in\mathcal{O}_{h}$. We remark that symmetric Euler elements are always contained in a subalgebra ${\mathfrak{h}}\simeq\mathop{{\mathfrak{sl}}}\nolimits_{2}({\mathbb{R}})$ of the Lie algebra ${\mathfrak{g}}.$ Non-symmetric Euler elements are contained in a subalgebra ${\mathfrak{h}}\simeq\mathop{{\mathfrak{gl}}}\nolimits_{2}({\mathbb{R}})$ of ${\mathfrak{g}}$ (cf. [MN22, Lem. 2.15]). Let $G$ be a Lie group, assume that $\tau$ is an Euler involution that integrates to an involution on $G$, then $G_{\tau}=G\rtimes_{\tau}\mathbb{Z}_{2}$ is well-defined and one has a continuous homomorphism $\epsilon\colon G_{\tau}\to\{\pm 1\}$ where
@@
so that $G\trianglelefteq G_{\tau}$ is a normal subgroup of index $2$ and $G\tau=G_{\tau}\setminus G$. Note that if $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ or $G$ is simply connected, then $\tau$ always integrate to an involution on $G$. For sake of simplicity assume here that that $G$ is connected and $G$ is center free.
@@
$\mathcal{G}_{E}{:=\mathcal{G}_{E}({G_{\tau}})}:={\{(h,\tau_{h})\in{\mathfrak{g}}\times G\tau\colon\,h\in\mathcal{E}({\mathfrak{g}})\}}$
is called the abstract Euler wedge space of ${G_{\tau}}$ and an element $(h,\tau_{h})\in\mathcal{G}_{E}$ is an Euler couple or Euler wedge. In the case $G$ is not center free or $\tau$ is not an Euler involution, Euler wedges are couples $(h,\sigma)\in{\mathfrak{g}}\times G\tau$ such that $h\in\mathcal{E}({\mathfrak{g}})$ and $\sigma$ is an involution such that $\mathop{{\rm Ad}}\nolimits\sigma=\mathop{{\rm Ad}}\nolimits\tau_{h}$. Given a fixed couple $W_{0}=(h,\tau_{h})\in\mathcal{G}_{E}$, the orbits
@@
Covariance of Euler wedges. A $G_{\tau}$-covariant action on $\mathcal{G}_{E}$ can be defined as follows. Consider the twisted adjoint action of $G_{\tau_{h}}$ which changes the sign on odd group elements:
@@
Order structure on Euler wedges. Given a “positive” cone in the Lie algebra one can define an order structure: Given an $\mathop{{\rm Ad}}\nolimits^{\epsilon}(G_{\tau_{h}})$-invariant pointed closed convex cone $C_{\mathfrak{g}}\subseteq{\mathfrak{g}}$, one defines the subsets
@@
([Ne22, Thm. 2.16]). Then the partial order on the orbit $G.W\subseteq\mathcal{G}_{E}$ is defined by
@@
In particular, $g.W\leq W$ is equivalent to $g\in S_{W}$. As an example, consider $G=\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})$ and its $\mathbb{Z}_{2}$-graded extension $G_{\tau_{h}}$. Then, if $C$ is chosen to be trivial, we are in the situation where there are no wedge inclusions, hence in the de Sitter wedge geometry. If we choose
@@
we are in the situation where there are wedge inclusions, hence of the circle $S^{1}$ wedge geometry picture. Namely the choice of $C$ influences the geometry of the model.
Locality and duality properties of Euler wedges. The notion of a “causal complement” is defined on the abstract wedge space as follows: For $W=(h,\tau)\in\mathcal{G}_{E}$, we define the dual wedge by $W^{\prime}:=(-h,\tau){=\tau.W}$. Note that $(W^{\prime})^{\prime}=W$ and $(gW)^{\prime}=gW^{\prime}$ for $g\in G$ by (3). The inclusion property of wedges fits with the duality property in the sense that the expression $W_{1}\subset W_{2}^{\prime}$ is meaningful. We shall refer to this condition as the locality property of the Euler couples. This relation aligns with the geometric interpretation within the context of wedge domains in spacetime manifolds.
Note that it is not always the case that $W_{0}^{\prime}\in\mathcal{W}_{+}(W_{0})$. As an example, consider the translation-dilation group $G$ acting on the real line. A wedge is a right half-line $I={\mathbb{R}}^{+}+a$ and is associated 1-1 to the positive dilations $\delta_{I}(t)x=e^{t}(x-a)+a$ where $x\in{\mathbb{R}}$. The complement interval $I^{\prime}$ is a negative half-line $I^{\prime}={\mathbb{R}}^{-}+a$, it has associated the one-parameter group $\delta_{I^{\prime}}(t)=\delta_{I}(-t)$. There exists no $g\in G$ such that $gI=I^{\prime}$ and one needs to include the $\tau_{I}$ to reflect $I$ to $I^{\prime}$, namely $I^{\prime}\in\mathcal{W}(I)$ but $I^{\prime}\not\in\mathcal{W}_{+}(I)$. This correspond to the fact that $h$ is not symmetric in ${\mathfrak{g}}$, see [MN21]. We further remark that if $Z(G)$ is non trivial, many abstract wedges correspond to an Euler element. This case is discussed in [MN21, Sect. 2.4.2].
We have now established a suitable framework for defining an abstract index set for Quantum Field Theory. Notably, this structure extends beyond well-known examples within the physics context. One can classify the Lie algebras ${\mathfrak{g}}$ supporting a three-grading, hence Euler wedges. In particular, the Lie group $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ and its extension $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})_{\tau_{h}}$ generalize the fundamental examples outlined in the previous section. If a simple real Lie algebra ${\mathfrak{g}}$ supports Euler elements then Theorem 3.10 in [MN21] specifies that its restricted root system must be one of the following:
@@
Here is the list of three graded Lie algebras:
@@ 2.3 Wedge domains in causal homogeneous spaces
Let $G$ be a connected Lie group, $H$ a closed subgroup. The associated homogeneous space is given by the quotient $M=G/H$. It is said to be causal if there exists a $G$-invariant field of pointed generating closed convex cones $\{C_{m}\}_{m\in M}$ with $C_{m}\subset T_{m}(M)$ the tangent space at the point $m\in M$. The Minkowski space, as well as the de Sitter space, is then a causal homogeneous space, because of the causal structure given by the Minkowksi metric.
@@
Motivated by the Bisognano–Wichmann property (BW) and its consequences in AQFT, the modular flow should be timelike future-oriented. It corresponds to the inner time evolution of Rindler wedges (see [CR94] and also [BB99, BMS01], [CLRR22, §3]). Then the connected component
@@
is the natural candidate for a wedge domain if $eH\in W_{M}^{+}(h)$ (or its boundary). As a concrete example, in Minkowski and de Sitter spacetimes the wedge $W_{\Lambda}$ is determined as the set of points $x\in M$ such that the vector field at $x$ associated to the one-parameter group $\Lambda$ of boosts is timelike and future pointing.
@@
Given a real connected simple Lie group $G$ such that the Lie algebra ${\mathfrak{g}}$ contains an Euler element $h\in\mathcal{E}({\mathfrak{g}})$, one can construct a non-compactly causal symmetric space as $M=G/H$ where $H$ is an open subgroup of $G^{\tau}$, the centralizer of the involution $\tau=\tau_{h}\theta$ with $\theta$ Cartan involution of ${\mathfrak{g}}$ such that $\theta(h)=-h$ that integrates on $G$. Note that within this picture $h\in T_{\textbf{e}}(M)\simeq{\mathfrak{g}}^{-\tau}$ where $\textbf{e}=eH$ is the basepoint of $M$. Furthermore one can define a $G$-invariant field of pointed generating closed convex cones $C_{\textbf{p}}\subset T_{\textbf{p}}(M)$ such that $h\in C_{\textbf{e}}^{\circ}$, hence $\textbf{e}\in W_{M}^{+}(h)$, see [MNO23a]. In this way one can associate a wedge domain $W(h)\subset M$ to any $h\in\mathcal{E}({\mathfrak{g}})$. A classification result of the local structure of such non compactly-causal homogenoeus spaces and connectedness of $W^{+}_{M}(h)$ is described in [MNO23a, Thm. 4.21], [MNO23b, Thm. 7.1].
As an example, consider the de Sitter space $\mathop{{\rm dS}}\nolimits$ of dimension $d$ and its symmetry group, the Lorentz group, $G=\mathop{{\rm SO}}\nolimits(1,d)_{e}$. Let $h:=h_{W^{\mathop{{\rm dS}}\nolimits}_{R}}$ be the generator of the one-parameter group of boosts in (1). Then following [MNO23a], one can consider the involution $\tau=\tau_{h}\theta$ on the Lie algebra ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(1,d)$ where $\tau_{h}$ is the Euler involution associated to the Euler element $h$ and $\theta$ is the Cartan involution of the Lie algebra ${\mathfrak{g}}$ promoted to $G$. In particular
@@
Let $\textbf{e}_{1}=(0,1,0,\ldots,0)\in\mathop{{\rm dS}}\nolimits\subset{\mathbb{R}}^{1,d}$ and $G_{\textbf{e}_{1}}$ be its stabilizer group, then the centralizer of $\tau$ in $G$ is described explicitly by
$G^{\tau}\simeq G_{\textbf{e}_{1}}\rtimes\{1,r_{1,2}(\pi)\}\simeq G_{\textbf{e}_{1}}\rtimes{\mathbb{Z}}_{2}$
where $r_{1,2}(\pi)$ is the $\pi$-rotation in the first two spatial coordinates. In particular $G^{\tau}$ has two connected components and $G_{\textbf{e}_{1}}$ is the identity component fixing $\textbf{e}_{1}$. Then de Sitter space $\mathop{{\rm dS}}\nolimits$ is isomorphic to $G/G_{\textbf{e}_{1}}$ which is a causal symmetric space but the quotient $G/G^{\tau}$ is the projective de Sitter space which is not a causal space: a global time orientation time orientation is missing. With the identification $\mathop{{\rm dS}}\nolimits\simeq G/G_{\textbf{e}_{1}}$ one has that $W_{M}^{+}(h)=W^{\mathop{{\rm dS}}\nolimits}_{R}$. We remark that what happens in this example is a general fact, namely when $h$ is a symmetric Euler element, then $M=G/G^{\tau}$ is never a causal manifold [MNO23a, Thm. 4.19].
@@ 2.4 The geometry of nets of real subspaces
In this section we recall some fundamental properties of the geometry of standard subspaces and one-particle nets. We refer to [LRT78, BGL02, MN21, NÓ17, Lo08] for more details. Particles in Quantum Field Theory are defined better through their symmetries rather than pointwise properties and their meaning is clear when the symmetry group is large enough. For instance, one-particle states of a free Quantum Field Theory on Minkowski spacetime are unit vectors of the Hilbert space supporting a unitary irreducible positive energy representation of the Poincaré group.
The free field is the most fundamental model in AQFT. We will see how the language of the standard subspaces captures the right features of the free model to allow a generalization of the AQFT on abstract sets of wedges and on causal symmetric spaces.

### v2

@@ 2.2 Euler elements and abstract wedges
We now introduce the natural general framework in which all these examples fit, clarifying the previously mentioned differences regarding the inclusion properties of wedge regions. The primary reference for this is [MN21].
@@
Then $\tau_{h}(y_{j})=\exp(i\pi\mathop{{\rm ad}}\nolimits h)(y_{j})=(-1)^{j}y_{j}$ for $y_{j}\in{\mathfrak{g}}_{j}(h)$ defines an involutive automorphism of ${\mathfrak{g}}$ called an Euler involution on ${\mathfrak{g}}$.
We denote $\mathcal{E}({\mathfrak{g}})$ the set of Euler elements in ${\mathfrak{g}}$. The orbit of an Euler element $h$ under the group $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})=\langle e^{\mathop{{\rm ad}}\nolimits{\mathfrak{g}}}\rangle$ of inner automorphisms is denoted with $\mathcal{O}_{h}=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})h\subseteq{\mathfrak{g}}$. We say that $h$ is symmetric if $-h\in\mathcal{O}_{h}$. In a simple Lie algebra ${\mathfrak{g}}$, the symmetric Euler elements are always contained in a subalgebra ${\mathfrak{h}}\simeq\mathop{{\mathfrak{sl}}}\nolimits_{2}({\mathbb{R}})$ and non-symmetric Euler elements are contained in a subalgebra ${\mathfrak{h}}\simeq\mathop{{\mathfrak{gl}}}\nolimits_{2}({\mathbb{R}})$ (see [MN22, Lem. 2.15] for a more general case). Let $G$ be a connected Lie group with Lie algebra ${\mathfrak{g}}$ and $\mathcal{E}({\mathfrak{g}})\neq\emptyset$, assume that $\tau$ is the involution on $G$ integrating an Euler involution on ${\mathfrak{g}}$. Then $G_{\tau}=G\rtimes_{\tau}\mathbb{Z}_{2}=G\rtimes\{e,\tau\}$ is well-defined and one has a continuous homomorphism $\epsilon\colon G_{\tau}\to\{\pm 1\}$ where
@@
so that $G\trianglelefteq G_{\tau}$ is a normal subgroup of index $2$ and $G\tau=G_{\tau}\setminus G$. Note that if $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ or $G$ is simply connected, then $\tau_{h}$ always integrate to an involution on $G$ when $h\in\mathcal{E}({\mathfrak{g}})$.
@@
$\mathcal{G}_{E}{:=\mathcal{G}_{E}({G_{\tau}})}:={\{(h,\sigma)\in{\mathfrak{g}}\times G\tau\colon\,h\in\mathcal{E}({\mathfrak{g}}),\mathop{{\rm Ad}}\nolimits\sigma=\tau_{h}\}}$
is called the abstract Euler wedge space of ${G_{\tau}}$, an element $(h,\sigma)\in\mathcal{G}_{E}$ is an Euler couple or Euler wedge and $\sigma$ is called Euler involution. For sake of simplicity we here assume here that that $G$ is center free. In this case every Euler element identifies an Euler couple $(h,\tau_{h})\in\mathcal{G}_{E}$ uniquely. One can also define more general abstract wedges without referring to Euler elements, see [MN21, Sect. 2].
Given a fixed couple $W_{0}=(h,\tau_{h})\in\mathcal{G}_{E}$, the orbits
@@
Covariance of Euler wedges. A $G_{\tau}$-covariant action on $\mathcal{G}_{E}$ can be defined as follows. Consider the twisted adjoint action of $G_{\tau}$ which changes the sign on odd group elements:
@@
Order structure on Euler wedges. Given an $\mathop{{\rm Ad}}\nolimits^{\epsilon}(G_{\tau})$-invariant pointed closed convex cone $C_{\mathfrak{g}}\subseteq{\mathfrak{g}}$, consider the subsets
@@
where $G_{W}:=\{g\in G\colon g.W=W\}$, see [Ne22, Thm. 2.16]. Then the partial order on the orbit $G.W\subseteq\mathcal{G}_{E}$ is defined by
@@
In particular, $g.W\leq W$ is equivalent to $g\in S_{W}$. As an example, consider $G=\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})$ and its $\mathbb{Z}_{2}$-graded extension $G_{\tau_{h}}$. If $C$ is chosen to be trivial, we arrive at a situation in which no non-trivial wedge inclusions exist, corresponding to the de Sitter wedge geometry. If we choose
@@
we are in the situation where there are proper wedge inclusions, hence the case of the circle $S^{1}$ wedge geometry. Namely the choice of $C$ influences the geometry of the model.
Duality and Locality properties of Euler wedges. For $W=(h,\tau_{h})\in\mathcal{G}_{E}$, we define the dual wedge by $W^{\prime}:=(-h,\tau_{h}){=\tau_{h}.W}$. This definition aligns with the geometric relation of a wedge with its causal complement in spacetime manifolds. Note that $(W^{\prime})^{\prime}=W$ and $(gW)^{\prime}=gW^{\prime}$ for $g\in G$ by (5). Since the wedge inclusions and the dual wedge are well defined notions on couples of Euler wedges we can state the locality property of Euler couples $W_{1}$ and $W_{2}$ that satisfy $W_{1}\subset W_{2}^{\prime}$.
Note that it is not always the case that $W_{0}^{\prime}\in\mathcal{W}_{+}(W_{0})$. As an example, consider the translation-dilation group $G$ acting on the real line. A wedge is a right half-line $I={\mathbb{R}}^{+}+a$ and is associated 1-1 to the positive dilations $\delta_{I}(t)x=\exp(h_{I}t)=e^{t}(x-a)+a$ where $x\in{\mathbb{R}}$. The complement interval $I^{\prime}$ is a negative half-line $I^{\prime}={\mathbb{R}}^{-}+a$, it has associated the one-parameter group $\delta_{I^{\prime}}(t)=\delta_{I}(-t)$. There exists no $g\in G$ such that $gI=I^{\prime}$ and one needs to include the $\tau_{I}$ to reflect $I$ to $I^{\prime}$, namely $I^{\prime}\in\mathcal{W}(I)$ but $I^{\prime}\not\in\mathcal{W}_{+}(I)$. This correspond to the fact that $h$ is not symmetric in ${\mathfrak{g}}$, see [MN21]. We further remark that if $Z(G)$ is non trivial, many abstract wedges correspond to an Euler element, and one can define a notion of “twisted complements” for a more general notion of complement wedges. In this case it is not always the case that $W^{\prime}_{0}\in\mathcal{W}_{+}(W_{0})$ even if $h$ is symmetric. This case is discussed in [MN21, Sect. 2.4.2].
We have now established a suitable framework for defining an abstract index set for Quantum Field Theory. Notably, this structure extends beyond well-known examples within the physics context. One can classify the real Lie algebras ${\mathfrak{g}}$ supporting a three-grading given by Euler elements. In particular, the Lie group $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$, hence its extension $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})_{\tau_{h}}$, supports Euler wedges. This generalizes the fundamental examples outlined in the previous section. By Theorem 3.10 in [MN21] if a simple real Lie algebra ${\mathfrak{g}}$ supports Euler elements then its restricted root system has to be one of the following
@@
The theorem also classify the number of inequivalent $\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$-orbits. Here is the list of three graded real Lie algebras:
@@ 2.3 Wedge domains in causal homogeneous spaces
Let $G$ be a connected Lie group, $H$ a closed subgroup. The associated homogeneous space is given by the quotient $M=G/H$. It is said to be causal if there exists a $G$-invariant field of pointed generating closed convex cones $\{C_{m}\}_{m\in M}$ with $C_{m}\subset T_{m}(M)$ the tangent space at the point $m\in M$. The Minkowski space, as well as the de Sitter space, is a causal homogeneous space, because of the causal structure given by the Minkowksi metric.
@@
Then the connected component
@@
is the natural candidate for a wedge domain if $eH\in W_{M}^{+}(h)$ (or its boundary). Motivated by the Bisognano–Wichmann property (BW) and its consequences in AQFT, the modular flow should be indeed timelike future-oriented in the wedge region. It corresponds to the inner time evolution of Rindler wedge (see [CR94] and also [BB99, BMS01], [CLRR22, §3]). As a concrete example, in Minkowski and de Sitter spacetimes the wedge $W_{\Lambda}$ is determined as the set of points $x\in M$ such that the vector field at $x$ associated to the one-parameter group $\Lambda$ of boosts is timelike and future pointing.
@@
Given a real connected simple Lie group $G$ such that the Lie algebra ${\mathfrak{g}}$ contains an Euler element $h\in\mathcal{E}({\mathfrak{g}})$, one can construct a non-compactly causal symmetric space as $M=G/H$ where $H$ is an open subgroup of $G^{\tau_{nc}}$, the centralizer of the involution $\tau_{nc}$ that integrates $\tau=\tau_{h}\theta$ on $G$ where $\theta$ is a Cartan involution of ${\mathfrak{g}}$ such that $\theta(h)=-h$. Note that within this picture $h\in T_{\textbf{e}}(M)\simeq{\mathfrak{g}}^{-\tau}$ where $\textbf{e}=eH$ is the basepoint of $M$. Furthermore one can define a $G$-invariant field of pointed generating closed convex cones $C_{\textbf{p}}\subset T_{\textbf{p}}(M)$ such that $h\in C_{\textbf{e}}^{\circ}$, hence $\textbf{e}\in W_{M}^{+}(h)$, and $C_{\textbf{e}}$ is $H$-invariant, see [MNO23a]. In this way one can associate a wedge domain $W(h)\subset M$ to any $h\in\mathcal{E}({\mathfrak{g}})$. A classification result of the local structure of such non compactly-causal homogeneous spaces and connectedness of $W^{+}_{M}(h)$ is described in [MNO23a, Thm. 4.21] and [MNO23b, Thm. 7.1].
As an example, consider the de Sitter space $\mathop{{\rm dS}}\nolimits$ of dimension $d$ and its symmetry group, the Lorentz group, $G=\mathop{{\rm SO}}\nolimits(1,d)_{e}$. Let $h:=h_{W^{\mathop{{\rm dS}}\nolimits}_{R}}$ be the generator of the one-parameter group of boosts in (3). Then following [MNO23a], one can consider the involution $\tau=\tau_{h}\theta$ on the Lie algebra ${\mathfrak{g}}=\mathop{{\mathfrak{so}}}\nolimits(1,d)$ where $\theta$ is the Cartan involution of the Lie algebra ${\mathfrak{g}}$ promoted to $G$. In particular
@@
Let $\textbf{e}_{1}=(0,1,0,\ldots,0)\in\mathop{{\rm dS}}\nolimits\subset{\mathbb{R}}^{1,d}$ and $G_{\textbf{e}_{1}}$ be its stabilizer group, then the centralizer of $\tau_{nc}$ in $G$ is described explicitly by
$G^{\tau_{nc}}\simeq G_{\textbf{e}_{1}}\rtimes\{1,r_{1,2}(\pi)\}\simeq G_{\textbf{e}_{1}}\rtimes{\mathbb{Z}}_{2}$
where $r_{1,2}(\pi)$ is the $\pi$-rotation in the first two spatial coordinates. In particular $G^{\tau_{nc}}$ has two connected components and $G_{\textbf{e}_{1}}$ is the identity component fixing $\textbf{e}_{1}$. Then de Sitter space $\mathop{{\rm dS}}\nolimits$ is isomorphic to $G/G_{\textbf{e}_{1}}$ which is a causal symmetric space but the quotient $G/G^{\tau_{nc}}$ is the projective de Sitter space which is not a causal space. With the identification $\mathop{{\rm dS}}\nolimits\simeq G/G_{\textbf{e}_{1}}$ one has that $W_{M}^{+}(h)=W^{\mathop{{\rm dS}}\nolimits}_{R}$. In general, let ${\mathfrak{g}}$ be a non compactly causal simple Lie algebra, $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ and $h\in\mathcal{E}({\mathfrak{g}})$, then $M=G/G^{\tau_{nc}}$ is a causal if and only if $h$ is not symmetric. Details are in [MNO23a, Thm. 4.19]).
@@ 2.4 The geometry of nets of real subspaces
In this section we recall some fundamental properties of the geometry of the standard subspaces and the one-particle nets. We refer to [LRT78, BGL02, MN21, NÓ17, Lo08] for more details. Particles in Quantum Field Theory are defined through their symmetries rather than pointwise properties and their meaning is clear when the symmetry group is large enough. For instance, one-particle states of a free Quantum Field Theory on Minkowski spacetime are unit vectors of the Hilbert space supporting a unitary irreducible positive energy representation of the Poincaré group.
The free field constitutes a fundamental model in AQFT. We will see how the language of the standard subspaces captures many of the features of the free models and how this allows a generalization of the AQFT framework on abstract sets of wedges and on causal symmetric spaces.